VAN  DER  WAALS  COMPLEXES  OF  3-HYDROXYTETRAHYDROFURAN:  3-

HYDROXYTETRAHYDROFURAN-H_{2}O AND 3-HYDROXYTETRAHYDROFURAN-Ar

\underline{RICHARD J. LAVRICH}, CHARLES R. TOROK, and MICHAEL J. TUBERGEN, Department of Chemistry, Kent State University, Kent, OH 44242.

Microwave spectra have been measured for the most abundant isotopic species of two van der Waals complexes of 3hydroxytetrahydrofuran: 3-hydroxytetrahydrofuran-H_{2}O and 3-hydroxytetrahydrofuran-Ar. For the H_{2}O complex the fitted rotational constants suggest that the water molecule forms a double hydrogen bond with the furan by donating a hydrogen bond to the ring oxygen and accepting a hydrogen bond from the hydroxyl group. Preliminary Stark effect experiments have been performed, yielding μ_{a} = 1.2 (3), μ_{b} = 1.82 (1), and μ_{c} = 0.7 (4) D. The rotational constants of the argon complex are fit best by a model that has the Ar atom situated on the same side of the ring as the hydroxyl group.